OpenStudy (anonymous):

Find the geometric mean of the pairs of numbers. 45 and 5.

jimthompson5910 (jim_thompson5910):

Hint: the geometric mean between x and y is \(\large \sqrt{x*y}\)

OpenStudy (anonymous):

is 45 x? and 5 y?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

\(\large \sqrt{x*y} = \sqrt{45*5} = ??\)

OpenStudy (anonymous):

33.54..

jimthompson5910 (jim_thompson5910):

hmm how did you get that?

OpenStudy (anonymous):

By doing exactly what you said..

jimthompson5910 (jim_thompson5910):

45 times 5 = ??

OpenStudy (anonymous):

I did sqrt?

jimthompson5910 (jim_thompson5910):

you do that first

jimthompson5910 (jim_thompson5910):

then you square root

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np